OpenStudy (anonymous):

a rectangular porch has a perimeter of 49 ft and a width of 6.5 ft. what is the area of the porch in square ft?

OpenStudy (mathstudent55):



OpenStudy (mathstudent55):

The perimeter is the sum of the two lengths and the two widths: x + x = 6.5 + 6.5 = 49 Solve this equation for x, the length. Then multiply the length and the width to find the area.

OpenStudy (anonymous):

the length is 18ft

OpenStudy (mathstudent55):

2x + 13 = 49 2x = 36 x = 18 Correct, the length is 18 ft. The area is: A = LW = 18 ft x 6.5 ft = 117 ft^2 The answer is: the area is 117 square feet.